OpenStudy (anonymous):

Resolve into Partial fractions xcube divided by 1-xcube

OpenStudy (anonymous):

right is it x^2 under is 1-x^2 ?

OpenStudy (anonymous):

Actually its X raise to power 3 by 1-Xraise to power 3

OpenStudy (anonymous):

\[\LARGE \frac{x^3}{(1-x)^3}\] this or \[\LARGE \frac{x^3}{1-x^3}\] this ?

OpenStudy (anonymous):

\[\LARGE \frac{x^3}{1-x^3}=\frac{x^3}{1^3-x^3}\] \[\LARGE \frac{x^3}{(1-x)(1^2+x\cdot 1+x^2)}=\frac{x^3}{(1-x)(1+x+x^2)}\] \[\LARGE \frac{x^2}{(1-x)}\cdot \frac{x}{(1+x+x^2)}=...\] go on...

OpenStudy (anonymous):

ok, Thanks :)